OpenStudy (anonymous):

1. 8/9 - 10/3 = a. -2/9 b. -1/9 c. -2 4/9 2. -4/5 - 5/6 = a. -1/30 b. -49/30 c. -9/30 3. 1/2 - (-5/9) = a. 6/9 b. 6/18 c. 19/18 4. -1.3 + -3.8 = a. 5.1 b. -5.1 c. 2.5 5. -2.87 + 4.2 = a. 1.33 b. -1.33 c. -7.07 6. 82.5 - -89.4 = a. -7 b. 7 c. 171.9 7. -5.89 - 6.1 = a. -0.21 b. 11.99 c. -11.99

OpenStudy (anonymous):

1) \[\frac{ 8 }{ 9 } -\frac{ 10 }{ 3 }(\frac{ 30 }{ 9 }) =-\frac{ 22 }{ 9 } or -2\frac{ 4 }{ 9 }\] 2) \[-\frac{ 4 }{ 5 }(\frac{ 24 }{ 30 })-\frac{ 5 }{ 6 }(\frac{ 25 }{ 30 })=-\frac{ 49 }{ 30 }\] 3) \[\frac{ 1 }{ 2 }(\frac{ 9 }{ 18 }) - (-\frac{ 5 }{ 9 })(\frac{ 10 }{ 18 }) = \frac{ 19 }{ 18 }\] 4) \[-1.3+-3.8=-5.1\] 5) \[-2.87 +4.2 = 1.33 or 4.2 -2.87 = 1.33\] 6) \[82.5 + 89.4 = 171.9\] 7) \[-5.89 - 6.1 = -11.99\]